# Daily Land Surface Temperature Reconstruction in Landsat Cross-Track Areas Using Deep Ensemble Learning With Uncertainty Quantification

### v1

@@ Abstract
Many real-world applications rely on land surface temperature (LST) data at high spatiotemporal resolution. In complex urban areas, LST exhibits significant variations, fluctuating dramatically within and across city blocks. Landsat provides high spatial resolution data at 100 meters but is limited by long revisit time, with cloud cover further disrupting data collection. Here, we propose DELAG, a deep ensemble learning method that integrates annual temperature cycles and Gaussian processes, to reconstruct Landsat LST in complex urban areas. Leveraging the cross-track characteristics and dual-satellite operation of Landsat since 2021, we further enhance data availability to 4 scenes every 16 days. We select New York City, London and Hong Kong from three different continents as study areas. Experiments show that DELAG successfully reconstructed LST in the three cities under clear-sky (RMSE = 0.73-0.96 K) and heavily-cloudy (RMSE = 0.84-1.62 K) situations, superior to existing methods. Additionally, DELAG can quantify uncertainty that enhances LST reconstruction reliability. We further tested the reconstructed LST to estimate near-surface air temperature, achieving results (RMSE = 1.48-2.11 K) comparable to those derived from clear-sky LST (RMSE = 1.63-2.02 K). The results demonstrate the successful reconstruction through DELAG and highlight the broader applications of LST reconstruction for estimating accurate air temperature. Our study thus provides a novel and practical method for Landsat LST reconstruction, particularly suited for complex urban areas within Landsat cross-track areas, taking one step toward addressing complex climate events at high spatiotemporal resolution.
@@ I Introduction
Continuous land surface temperature (LST) data from satellites is fundamental to achieve sustainable development, to understand climate, and to analyze climate impacts on planetary health [1]. Many of the real-world applications rely on satellite LST data, including monitoring wildfires, weather forecasting, and understanding temperature-related mortality [2, 3, 4]. Although satellites can provide LST data on a large scale, these data are often incomplete due to cloudy weather conditions and sometimes sensor failure [5, 6, 7]. For high spatial resolution satellite data, due to sensor technology, the trade-off between spatial and temporal resolutions further restricts daily data collection. Although daily LST data is crucial for real-world applications, currently we don’t have good enough daily LST data at high spatiotemporal resolution [8].
@@
Figure 1: Spatial resolution comparison of three types of satellite-derived temperature products: gridMET (4 km), MODIS (1 km), Landsat (30 m resampled from 100 m). The overlaid polygons are the New York City merged zip-code boundaries that have daily counts of emergency department visit data. But at the spatial resolution of gridMET (4 km) and MODIS (1 km), these existing LST data are not sufficient to capture temperature variations. Landsat LST can capture temperature variations but is limited by temporal frequency and clouds.
At the kilometer resolution of MODIS and other commonly used LST data, it is impossible to capture temperature heterogeneity within a city [10]. In complex urban areas, temperature changes dramatically over short distances, ranging from a few meters to street blocks [20]. For example, in New York City as shown in Figure 1, many urban parks have been able to reduce daytime temperature with vegetation. The city has complex street blocks exhibiting large temperature variations, but these variations are not captured from the coarse LST data, such as the 4 km resolution gridMET and the 1 km resolution MODIS (Figure 1). The coarse-resolution LST data and the resulting coarse-resolution water masks also make data unavailable in many coastal urban areas with high population density such as Manhattan (Figure 1). Having fine-scale LST data is therefore critical due to the complex landscape, vast temperature heterogeneity, and high population concentration in cities that are more vulnerable to climate impacts [21]. Additionally, in real-world applications, many datasets are aggregated to small geographical units, such as zip codes (Figure 1). Many socioeconomic and health datasets are available at this level to support research on various topics, including air pollution, respiratory diseases, and temperature-related mortality [22, 23, 24]. However, the development of high resolution temperature data lacks behind, and many real-world applications had to settle to the coarse-resolution LST data [25, 26, 27].
@@
Landsat is the only widely-used satellite series with sufficiently high spatial resolution to capture temperature heterogeneity within a city. Landsat has a revisit time of 16 days, an often-used excuse (in addition to clouds) to eliminate its potential to generate daily LST data. In the existing literature, only a very small subset has explored Landsat LST reconstruction. For example, Zhu et al. (2022) [28] proposed an ATC-based novel method to reconstruct cloudy LST based on a series of procedures including cloud removal, shadow identification, and grouping from spatially adjacent similar pixels. An earlier work used temporal Gaussian processes for daily Landsat LST reconstruction without considering spatial correlation [19]. As a restriction of Landsat’s long revisit time, both studies adopted an inter-annual approach to ensure sufficient data support, making the interpretation of ATC parameters that are supposed to have physical meaning difficult. These methods are also limited to dates with Landsat observations; on many dates without Landsat observations, the uncertainty drawn from these approaches can be large, but the reliability and uncertainty of the reconstruction are unknown.
@@
To generate high resolution LST data on a daily basis, especially in cities, we here propose a framework that consists of two parts: from the perspectives of both data and method. From the data perspective, we leverage the Landsat cross-track areas and highlight its overlooked role in LST reconstruction. Although Landsat has a revisit time of 16 days, the Landsat cross-track areas (2 observations per 16 days) cover many cities that can benefit from this special setting. For example, the densely-populated New York City is completely situated within two Landsat tracks (Figure 2). Two other cities–London and Hong Kong–in this study also have their majority within the cross-track areas. With both Landsat 8 and 9 satellites in operation since 2021, the frequency is further enhanced to 4 scenes per 16 days, making daily Landsat LST reconstruction for practical applications toward feasible. From the method perspective, in this study, we propose DELAG, a deep ensemble learning method incorporating enhanced ATC models and Gaussian processes for Landsat LST reconstruction. DELAG utilizes deep ensembles to achieve prediction intervals – first of its kind. The enhanced ATC model and Gaussian processes are used to link clear-sky pixels to cloud-covered pixels. There is no LST site within the study areas, and given that many real-world applications require air temperature (instead of LST) derived from satellites [29] and the ability of LST to generate accurate air temperature [30], we validate the proposed method by comparing the air temperature generated from clear-sky LST and reconstructed LST using a simple linear statistical method, a validation approach commonly used in the literature [31, 18].
@@
In the remainder of this paper, we first introduce and visualize the Landsat cross-track areas in Section II. We show that Landsat cross-track areas are sufficiently generalized, in many countries and regions, especially in mid- to high-latitude densely populated areas. We select three representative cities–New York City, London and Hong Kong–across three different continents as study areas here to demonstrate the generalization of the proposed framework. We then detail the proposed method (DELAG), the auxiliary data and validation strategies in Section IV. In Section V, we show the experimental results in the three cities, the ablation analysis, and the indirect validation (and application) with near-surface air temperature data. We discuss the broader impacts and conclude this study in Section VI.
II Landsat Cross-Track Areas
@@
The Landsat satellites are a series of polar-orbiting satellites that are designed for land use and land cover monitoring [32]. Since Landsat 4 launched in 1982, the Landsat satellites have thermal sensors onboard that can record land surface temperature. The current operating satellites (Landsat 8 and 9) have two thermal bands at 10.6-11.2 µm (Band 10) and 11.5-12.5 µm (Band 11). As a result of their polar-orbiting nature, each Landsat satellite has a revisit time of 16 days. Between two adjacent Landsat tracks (recording data on two different dates), there is an overlapped area that has dual observations [33]. The cross-track characteristics is often overlooked, as these areas are often assumed insignificant. But in fact, they cover a rather large area, especially at mid- to high-altitude regions, including some densely populated urban areas. Additionally, from a data perspective, we have the potential to recover the entire LST surface on Earth from the partial observations.
@@
In the following calculation and visualization, we demonstrate that the Landsat cross-track areas are not as insignificant as previously assumed. In the current specification, Landsat captures each scene in a dimension of 185 km $\times$ 180 km following its polar-orbiting trajectory [32]. In Figure 2, we show two sample scenes and highlight the cross-track areas that completely cover New York City. Landsat scenes have a width of 185 km ($L_{landsat}$) in a roughly horizontal direction, with a 8.2 shift parallel to latitude, leading to a horizontal coverage of 183.1 km. Every 16 days, a Landsat satellite travels 233 tracks, and then revisits. The ratio $\phi_{lat}$ between the horizontal length of all Landsat scenes within a revisit period (16 days, 233 tracks) and the latitude (with a length of $L_{lat}$) can be calculated as
@@
The ratio reaches its minimum (1.07) at the Equator, meaning that about 7% of the areas at the Equator are covered by two Landsat scenes per 16 days. The percentage of overlapping areas increases with latitude increasing and reaches to a ratio of 1.50 at 45 latitude: half of the areas at 45N and 45S are covered by two Landsat scenes per 16 days. We compute and visualize these Landsat cross-track areas in Figure 3, and the percentage of the cross-track areas in each country and region in Figure 4. Eleven countries and regions are 100% located in the Landsat cross-track areas, and many countries and regions have half of their areas located in the Landsat cross-track areas. Mid-to-high-latitude countries and regions, including most in Europe, Canada, and Russia have a high percentage of cross-track areas. At the metropolitan level, the cross-track areas cover many densely-populated urban areas including New York City (100%), London ($>$95% areas), and Hong Kong ($>$90% urban areas). As a combined result of cross tracks and dual satellites, we have four scenes per 16 days in these cities, making it toward feasible for daily Landsat LST reconstruction. In this study, we select New York City, London and Hong Kong as the three representative cities as study areas to demonstrate the generalization of the proposed framework and the proposed method.
Figure 3: Global areas covered by two or more Landsat scenes within a revisit period of 16 days, marked in red.
@@
III Study Areas and Data
III-A Study Areas
III-A1 New York City
New York City (NYC) is a megacity with the highest population density in the United States and is the primary case in this study for three reasons. First, the city’s health department releases daily syndromic surveillance data (e.g., respiratory emergency department visit) at the zip code level to the public. This dataset has been commonly used in real-world applications to investigate how various air pollutants and other environmental exposures affect human health [24, 34, 35]. However, currently, there is no daily LST data available that is able to distinguish temperature exposures at each zip code, making it impossible to address the impacts and disparities of temperature exposure within the city. Second, NYC is completely situated within the Landsat cross-track areas (Figure 2), making it an ideal study case for the proposed framework. Third, NYC has complex urban landscape and high temperature variations. With 8.3 million residents, the city has a population density of 11,000 per km2, about 110 residents per Landsat thermal pixel (100 m $\times$ 100 m). Given the complex urban setting, LST varies from locations to locations. High spatiotempral resolution LST data are therefore necessary.
III-A2 London

### v2

@@ Abstract
Many real-world applications rely on land surface temperature (LST) data at high spatiotemporal resolution. In complex urban areas, LST exhibits significant variations, fluctuating dramatically within and across city blocks. Landsat provides high spatial resolution data at 100 meters but is limited by long revisit time, with cloud cover further disrupting data collection. Here, we propose DELAG, a deep ensemble learning method that integrates annual temperature cycles and Gaussian processes, to reconstruct Landsat LST in complex urban areas. Leveraging the cross-track characteristics and dual-satellite operation of Landsat since 2021, we further enhance data availability to 4 scenes every 16 days. We select New York City, London and Hong Kong from three different continents as study areas. Experiments show that DELAG successfully reconstructed LST in the three cities under clear-sky (RMSE = 0.73-0.96 K) and heavily-cloudy (RMSE = 0.84-1.62 K) situations, superior to existing methods. Additionally, DELAG can quantify uncertainty that enhances LST reconstruction reliability. We further tested the reconstructed LST to estimate near-surface air temperature, achieving results (RMSE = 1.48-2.11 K) comparable to those derived from clear-sky LST (RMSE = 1.63-2.02 K). The results demonstrate the successful reconstruction through DELAG and highlight the broader applications of LST reconstruction for estimating accurate air temperature. Our study thus provides a novel and practical method for Landsat LST reconstruction, particularly suited for complex urban areas within Landsat cross-track areas, taking one step toward addressing complex climate events at high spatiotemporal resolution. Code and data are available at https://skrisliu.com/delag.
@@
This is the preprint version. For the final version, go to https://doi.org/10.1109/tgrs.2025.3643985.
@@ I Introduction
Continuous land surface temperature (LST) data from satellites is fundamental to achieve sustainable development, to understand climate, and to analyze climate impacts on planetary health [1]. Many of the real-world applications rely on satellite LST data, including monitoring wildfires, weather forecasting, and understanding temperature-related mortality [2, 3, 4]. Although satellites can provide LST data on a large scale, these data are often incomplete due to cloudy weather conditions and sometimes sensor failure [5, 6, 7]. For high spatial resolution satellite data, due to sensor technology, the trade-off between spatial and temporal resolutions further restricts daily data collection. Although daily LST data is crucial for real-world applications, currently, we don’t have good enough daily LST data at high spatiotemporal resolution [8].
Figure 1: Spatial resolution comparison of satellite-derived land surface temperature products in New York City: MODIS (1 km), Landsat (30 m resampled from 100 m). The overlaid polygons are the New York City merged zip-code boundaries that have daily counts of emergency department visit data. The coarse-resolution LST data is not sufficient to capture spatial variations and urban heterogeneity. Landsat LST can capture temperature variations but is limited by temporal frequency and clouds.
The most commonly used temperature data today is the MODIS observational, gridMET hybrid and ERA5 reanalysis data, with 1 km, 4 km and 11 km spatial resolution, respectively [9, 10]. But, at their kilometer resolution, it is impossible to capture temperature heterogeneity within a city [11]. In complex urban areas, temperature changes dramatically over short distances, ranging from a few meters to street blocks [12]. For example, in New York City as shown in Fig. 1, many urban parks have been able to reduce daytime temperature with vegetation. The city has complex street blocks exhibiting large temperature variations, but these variations are not captured from the 1 km resolution MODIS data (Fig. 1). The coarse-resolution LST data and the resulting coarse-resolution water masks also make data unavailable in many coastal urban areas with high population density such as Manhattan (Fig. 1). Having fine-scale LST data is therefore critical due to the complex landscape, vast temperature heterogeneity, and high population concentration in cities that are more vulnerable to climate impacts [13]. Additionally, in real-world applications, many datasets are aggregated to small geographical units, such as zip codes (Fig. 1). Many socioeconomic and health datasets are available at this level to support research on various topics, including air pollution, respiratory diseases, and temperature-related mortality [14, 15, 16]. However, the development of high resolution temperature data lacks behind, and many real-world applications had to settle to the coarse-resolution LST data [17, 18, 19].
@@
Landsat is the only widely-used satellite series with sufficiently high spatial resolution to capture temperature heterogeneity within a city. Landsat has a revisit time of 16 days, an often-used excuse (in addition to clouds) to eliminate its potential to generate daily LST data. In the existing literature, only a very small subset has explored Landsat LST reconstruction. For example, Zhu et al. [30] proposed an ATC-based novel method to reconstruct cloudy LST based on a series of procedures including cloud removal, shadow identification, and grouping from spatially adjacent similar pixels. An earlier work used temporal Gaussian processes for daily Landsat LST reconstruction without considering spatial correlation [29]. As a restriction of Landsat’s long revisit time, both studies adopted an inter-annual approach to ensure sufficient data support, making the interpretation of ATC parameters that are supposed to have physical meaning difficult. These methods are also limited to dates with Landsat observations; on many dates without Landsat observations, the uncertainty drawn from these approaches can be large, but the reliability and uncertainty of the reconstruction are unknown.
@@
To generate high resolution LST data on a daily basis, especially in cities, we here propose a framework that consists of two parts: from the perspectives of both data and method. From the data perspective, we leverage the Landsat cross-track areas and highlight their overlooked role in LST reconstruction. Although Landsat has a revisit time of 16 days, the Landsat cross-track areas (2 observations per 16 days) cover many cities that can benefit from this special setting. For example, the densely-populated New York City is completely situated within two Landsat tracks (Fig. 2). Two other cities—London and Hong Kong—in this study also have their majority within the cross-track areas. With both Landsat 8 and 9 satellites in operation since 2021, the frequency is further enhanced to 4 scenes per 16 days, making daily Landsat LST reconstruction for practical applications toward feasible. From the method perspective, in this study, we propose DELAG, a deep ensemble learning method incorporating enhanced ATC models and Gaussian processes for Landsat LST reconstruction. DELAG utilizes deep ensembles to achieve prediction intervals—the first of its kind. The enhanced ATC model and Gaussian processes are used to link observed pixels to cloud-covered pixels. There is no LST site within the study areas, and given that many real-world applications require air temperature (instead of LST) derived from satellites [31] and the ability of LST to generate accurate air temperature [32], we validate the proposed method by comparing the air temperature generated from observed LST and reconstructed LST using a simple linear statistical method, a validation approach commonly used in the literature [28, 33].
@@
In the remainder of this paper, we first provide a concise literature review on the current LST reconstruction methods in Section II. We then introduce and visualize the Landsat cross-track areas in Section III. We show that Landsat cross-track areas are sufficiently generalized, in many countries and regions, especially in mid- to high-latitude densely populated areas. We select three representative cities—New York City, London and Hong Kong—across three different continents as study areas here to demonstrate the generalization of the proposed framework. We then detail the proposed method (DELAG), the auxiliary data and validation strategies in Section V. In Section VI, we show the experimental results in the three cities, the ablation analysis, and the indirect validation (and application) with near-surface air temperature data. We discuss the broader impacts and conclude this study in Section VII.
II Literature on LST Reconstruction Methods
Reconstruction of LST data, from an application perspective, can be broadly categorized into two types: all-weather (at least daily) reconstruction and non-all-weather reconstruction. Based on this distinction, all-weather approaches can be further classified according to the sources or channels through which they obtain information about unobserved LST. These methods include spatial and/or temporal interpolation and their variants, data fusion integrating multiple satellites, passive microwave and ground-based observations, surface energy balance modeling, and numerical modeling of Earth system processes. Two comprehensive reviews of LST reconstruction methods are available in the literature [11, 34]. In this section, we primarily provide a focused overview relevant to daily LST reconstruction, with particular attention to ATC models, surface energy balance, and deep learning approaches. The proposed method can be understood as a hybrid approach that integrates information derived from spatiotemporal interpolation via ATC and Gaussian processes and surface energy balance through the output of numerical modeling (ERA5).
II-A Spatial and Temporal Interpolation With Land Surface Information
Spatial interpolation aims to gap-fill LST data using valid observations available at the same time point. In its narrow definition, this class of methods explicitly models spatial proximity—typically Euclidean distance—between observed and unobserved pixels, often employing techniques such as Gaussian processes and graph-based models [35]. More broadly, the notion of distance is extended to kernel-based or feature-based distances (also referred to as latent distances), where similarity is measured through spectral reflectance including vegetation index or latent representations of land surface characteristics [30, 36, 37]. Some studies further shift the problem from interpolation to a regression framework, where the objective becomes learning a mapping from input features $x$ to LST $T_{s}=f(x)$, often using machine learning models [22, 28]. By redefining distance in terms of feature similarity rather than geographic location, these models relax the constraint of spatial adjacency and are more adaptable to real-world conditions, where cloud cover is typically non-random and frequently occludes large contiguous regions.
Regardless of whether spatial interpolation is defined narrowly or broadly, its effectiveness depends on the availability of valid observations. On heavily overcast or rainy days, zero valid observations are not uncommon, and this challenge is compounded by the infrequent revisit times of many satellite sensors. Consequently, spatial interpolation methods are often applied to MODIS data, which provides partial daily observations [22, 38, 39]. To address data gaps more effectively, temporal interpolation—typically based on the ATC model—is frequently integrated prior to spatial methods [30, 40, 41, 42]. Still, these combined approaches are limited on days without any available observations.
II-B Data Fusion and Super-Resolution
Integrating observations from multiple data sources is a natural step toward improving data availability and reliability [43, 44, 45]. With the growing interest in data fusion and super-resolution techniques in machine learning, these methods have become increasingly popular for LST reconstruction [44, 46, 47]. In a typical machine learning framework, given a pair of low- and high-resolution data, the goal is to find a function $f(\cdot)$ to map low-resolution data to high resolution, i.e., $x_{\text{high}}=f(x_{\text{low}})$. For example, several spatiotemporal fusion methods have been applied to downscale GOES-16 LST data from 2 km to Landsat’s 100 m resolution [47]. Data fusion requires cloud-free pairs of low- and high-resolution images, which can be a limiting factor in real-world applications. To address this, Li et al. [48] first filled gaps in the low-resolution image using a clear reference image before applying a super-resolution neural network. While data fusion and super-resolution methods hold strong potential due to their data-driven nature, they remain limited by weather conditions and typically require more intensive data collection and training resources compared to other approaches. In some cases, low-resolution images may not be available at all. Microwave observations, which are not affected by cloud cover, offer some promise for obtaining all-weather LST data, but their current temporal frequency is still insufficient for daily reconstruction [49].
II-C Surface Energy Balance-Based Approaches
As the availability of observational data fundamentally limits methods based on interpolation and data fusion, a growing direction is to supplement these approaches with information derived from physical principles, particularly surface energy balance [50]. This additional step can take relatively simple forms, such as applying a linear correction to adjust for cloud effects [50], or comparing cloud-contaminated pixels with clear-sky reference images using similar-pixel techniques [51]. However, due to the complexity of heat transfer processes and surface energy balance calculations at fine spatial scales, such modeling is often oversimplified, and the representation of cloud effects within these calculations remains insufficient [34]. Numerical models and climate reanalysis data, which explicitly incorporate surface energy balance into calculations at the global scale, produce outputs that implicitly encode this physical information. Building on this foundation, recent studies have begun integrating numerical model outputs to estimate the difference between observed and cloud-covered LST as simulated by ERA5 [52]. These approaches involve first estimating the clear-sky and cloud-cover LST difference using ERA5 data, then applying the result to higher-resolution observational data for reconstruction [52, 53, 54].
II-D Linking Prior Work to the Proposed Method
The proposed method is primarily a hybrid approach that combines spatiotemporal interpolation (via ATC and Gaussian Processes) with information from surface energy balance, incorporated as a linear term based on coarse-resolution ERA5 data that is the output of numerical models. It differs from prior work by integrating these components into a single equation and treating ERA5 not as a spatially resolved image but as the areal mean, an assumption that, when combined with observational data from Landsat, enables reconstruction at the Landsat scale. Additional contributions of the proposed method include uncertainty quantification through deep ensemble learning, as well as a broader framework for leveraging Landsat cross-track areas to increase data frequency. Finally, we introduce a validation strategy based on near-surface air temperature data. While not entirely new, this approach is less commonly used than traditional in-situ LST validation that is infeasible in major metropolitan areas like the three major cities in this study.
III Landsat Cross-Track Areas
@@
The Landsat satellites are a series of polar-orbiting satellites that are designed for land use and land cover monitoring [55]. Since Landsat 4 launched in 1982, the Landsat satellites have thermal sensors onboard that can record land surface temperature. The current operating satellites (Landsat 8 and 9) have two thermal bands at 10.6–11.2 µm (Band 10) and 11.5–12.5 µm (Band 11). Because of polar orbiting, each Landsat satellite has a revisit time of 16 days. Between two adjacent Landsat tracks (recording data on two different dates), there is an overlapped area that has dual observations [56]. The cross-track characteristics are often overlooked, as these areas are commonly assumed to be insignificant. However, they actually cover a considerable area, especially in mid- to high-latitude regions, including some densely populated urban areas. Additionally, from a data perspective, we have the potential to recover the entire LST surface on Earth from the partial observations.
Figure 3: Global areas covered by two or more Landsat scenes (in red) within a revisit period of 16 days
@@
In the following calculation and visualization, we demonstrate that the Landsat cross-track areas are not as insignificant as previously assumed. In the current specification, Landsat captures each scene in a dimension of 185 km $\times$ 180 km following its polar-orbiting trajectory [55]. In Fig. 2, we show two sample scenes and highlight the cross-track areas that completely cover New York City. Landsat scenes have a width of 185 km ($L_{landsat}$) in a roughly horizontal direction, with a 8.2 shift parallel to latitude, leading to a horizontal coverage of 183.1 km. Every 16 days, a Landsat satellite travels 233 tracks, and then revisits. The ratio $\phi_{lat}$ between the horizontal length of all Landsat scenes within a revisit period (16 days, 233 tracks) and the latitude (with a length of $L_{lat}$) can be calculated as
@@
The ratio reaches its minimum (1.07) at the Equator, about 7% of the areas at the Equator are covered by two Landsat scenes per 16 days. The percentage of overlapping areas increases with latitude increasing and reaches to a ratio of 1.50 at 45 latitude: half of the areas at 45N and 45S are covered by two Landsat scenes per 16 days. We compute and visualize these Landsat cross-track areas in Fig. 3, and the percentage of the cross-track areas in each country and region in Fig. 4. Eleven countries and regions are 100% located in the Landsat cross-track areas, and many countries and regions have half of their areas located in the Landsat cross-track areas. Mid-to-high-latitude countries and regions, including most in Europe, Canada, and Russia have a high percentage of cross-track areas. At the metropolitan level, the cross-track areas cover many densely-populated urban areas including New York City (100%), London (>>95% areas), and Hong Kong (>>90% urban areas). As a combined result of cross tracks and dual satellites, we have four scenes per 16 days in these cities, making it toward feasible for daily Landsat LST reconstruction. In this study, we select New York City, London and Hong Kong as the three representative cities as study areas to demonstrate the generalization of the proposed framework and method.
@@
IV Study Areas and Data
IV-A Study Areas
IV-A1 New York City
New York City (NYC), located at 40.7N and 74.0W and fully covered by two Landsat scenes (Path 014 Row 032 and Path 013 Row 032), is the most densely populated megacity in the U.S. and serves as the primary case in this study for three reasons. First, NYC’s health department publicly releases daily syndromic surveillance data (e.g., respiratory emergency visits) at the zip code level. This dataset has been widely used to study the health impacts of air pollutants and other environmental exposures [16, 57, 58]. However, no daily LST data is currently available at the zip code scale, limiting our ability to assess temperature-related impacts and disparities within the city. Second, NYC lies entirely within Landsat’s cross-track coverage (Fig. 2), making it an ideal study area. Within the first 16-day circle (Jan 4-19) in 2023, Landsat 8 obtained data on the 4th and 13th, and Landsat 9 obtained data on the 5th and 12th, and then the circle repeats every 16 days, totaling 92 scenes in 2023. Third, NYC’s complex urban landscape and high temperature variability require fine-resolution data. With 8.3 million residents and a density of 11,000 per km2, about 110 people per 100 m $\times$ 100 m Landsat thermal pixel, LST varies substantially across the city. NYC has a humid subtropical climate, with its surroundings located in the transition zone between humid subtropical and humid continental climates. The daily mean temperature is about 0.7C in January and 25.3C in July. The typical daily minimum and maximum range is between -2.3 and 29.4C.
IV-A2 London